# On the maximal correlation of some stochastic processes

### v2

@@ Abstract
We consider the maximal correlation coefficient $R(X,Y)$ between two stochastic processes $X$ and $Y$. When $(X,Y)$ is a random walk, the result is a consequence of Csáki-Fischer identity and lower-semi continuity of $\text{Law}(X,Y)\to R(X,Y)$. When $(X,Y)$ are two-dimensional Lévy processes, we give an expression of $R(X,Y)$ via the covariance $\Sigma$ and the Lévy measure $\nu$ appeared in the Lévy-Khinchine formula. As a consequence, for two-dimensional $\alpha$-stable random variables $(X,Y)$ with $20<\alpha<$ 2, we give an expression of $R(X,Y)$ via the stability index $\alpha$ and the spectral measure $\tau$ of the $\alpha$-stable distribution. We also prove analogs and generalizations of Dembo-Kagan-Shepp-Yu inequality and Madiman-Barron inequality. Roughly speaking, we investigate the maximal correlation coefficient between two random selected “subvectors” $Y$ and $Z$ of a common random vector $X$. Besides, by using above new results, we recover several classical results.
@@ 1 Introduction
For two non-degenerate real random variables $X$ and $Y$, the Pearson correlation coefficient $\rho(X,Y)$ is defined by
@@
It measures the linear dependence of $X$ and $Y$. It is well-known that $\rho(X,Y)=0$ if $X$ and $Y$ are independent. However, the converse is not true in general. For two non-degenerate random variables $X$ and $Y$, the maximal correlation coefficient between $X$ and $Y$, introduced by Gebelein [Geb41], is given by
@@
where the supremum is taken for all measurable functions $\varphi$ and $\psi$ such that
$0<E[(\varphi(X))^{2}]<\infty,0<E[(\psi(Y))^{2}]<\infty.$
If $X$ or $Y$ is degenerate, we define $R(X,Y)=0$. The quantity $R(X,Y)$ measures the dependence between $X$ and $Y$. Note that $R(X,Y)=0$ if and only if $X$ and $Y$ are independent, see e.g. [Sar58].
It is well known that the maximal correlation coefficient has the following alternative definition: For a random variable $X$, denote by $L^{2}_{0}(X)$ the Hilbert space of real random variables $Z$ measurable with respect to the sigma-field $\sigma(X)$ generated by $X$ such that $E(Z)=0$ and $E(Z^{2})<\infty$. Then, $R(X,Y)$ equals the operator norm of the conditional expectation $\varphi(X)\mapsto E(\varphi(X)|Y)$ from $L^{2}_{0}(X)$ to $L^{2}_{0}(Y)$. Equivalently, we have that
$R^{2}(X,Y)=\sup\{E[(E(\varphi(X)|Y))^{2}]:E[\varphi(X)]=0,E[\varphi^{2}(X)]=1\}.$ (2)
Recently, the maximal correlation coefficient is generalized and studied by Dadoun and Youssef [DY21] in the context of free probability. But we will concentrate on classical notion of maximal correlation in the present paper.
The maximal correlation coefficient plays an important role in various areas of probability and statistics, including information theory [Cou16, MB07], hyper-contractivity ribbon of a pair of random variables and impossibility of non-interactive simulations of joint distributions [AG76, KA12, KA16], optimal transformation for regression [BF85], the spectral gap of Markov chains and the convergence theory of data augmentation algorithms [LWK94].
As pointed out by Rényi in [R5́9], it is often difficult to calculate the maximal correlation coefficient although it may seem to be not so difficult at the first sight. One reason is that the supremum in (1) could not be replaced by maximum in general. The explicit formula of the maximal correlation coefficient $R(X,Y)$ is known in a few cases. For example, when $(X,Y)$ is jointly Gaussian,
@@
see [CF60, Example 5 in Section 2]. They also found the maximal correlation for multinominal distributions and multihypergeometric distributions in [CF63, Section 6]. Bücher and Staud [BS25] have obtained the maximal correlation coefficient for bivariate Marshall Olkin exponential distribution. There are also several results for maximal correlation coefficients by using orthogonal polynomials, see [LBCnM06, LBSMn14, LBSMn98, Nev92, PX13, SM85, Ter83]. Among these results, [SM85] gives the maximal correlation coefficients for Dirichlet distributions, which provides sharp upper bounds for the maximal correlation coefficients for order statistics studied in [SM85, Ter83]. In [LBSMn98, Nev92], they studied the maximal correlation between the $i$-th and $j$-th records. In [PX13], they provided a unified approach for obtaining the maximal correlation coefficient for a subclass of Lancaster distributions. For maximal correlation coefficients for general distributions, we are aware of the following results. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are independent. Then, we have that
@@
This is known as Csáki-Fischer identity [CF63, Theorem 6.2], see also [Wit75, Theorem 1]. If $X_{1},X_{2},\ldots,X_{n}$ are independent and identically distributed (i.i.d.) non-degenerate random variables, then for $m\leq n$, we have that
@@
where $S_{k}=\sum_{i=1}^{k}X_{i}$ is the partial sum of $(X_{i})_{i}$, see [BDK05, DKS01, Nov04]. Let’s call the upper bound $R(S_{n},S_{m})\leq\sqrt{m/n}$ the Dembo-Kagan-Shepp (DKS) inequality. Yu [Yu08, Theorem 4.1] further generalized (4) to
@@
We are interested in the maximal correlation coefficient between two stochastic processes, e.g. two random walks, two Lévy processes or two random sub-vectors of a third common random vector.
To be more precise, let $(\xi_{n},\eta_{n})_{n\geq 1}$ be i.i.d. random vectors. Let $(S_{n},T_{n})_{n\geq 0}$ be a random walk starting from 00 with increments $(\xi_{n},\eta_{n})$, i.e.
@@
Let $S=(S_{n})_{n\geq 0}$ and $T=(T_{n})_{n\geq 0}$. Then, we have the following result:
Theorem 1.1 (Csáki-Fischer).
@@
The first equality is a direct consequence of Csáki-Fischer identity (Theorem 2.2) since
@@
For the second equality, we use the lower semi-continuity of $\text{Law}(X,Y)\mapsto R(X,Y)$ (Lemma 2.3). For self-containedness, we also provide a new probabilistic proof of the Csáki-Fischer identity. As a consequence of Theorem 1.1, the central limit theorem and the lower semi-continuity of $\text{Law}(X,Y)\mapsto R(X,Y)$ (Lemma 2.3), we give a new proof of (3). Similarly, by using the law of rare events instead of the central limit theorem, we see that (3) holds for bivariate Poisson distributions.
Next, we move to a new related problem. We have considered the maximal correlation coefficient for random walks. As Lévy processes are continuous analogs of random walks, it is natural to consider the maximal correlation coefficient between two correlated Lévy processes $(X_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$. We assume that $(X_{t},Y_{t})$ is a two dimensional Lévy process. We refer to Applebaum’s book [App09] for properties of Lévy processes. The important Lévy-Khinchine formula is well-known. In the two-dimensional case, it has the following form:
$E[e^{i(u_{1}X_{t}+u_{2}Y_{t})}$ $]=\exp\left(t\left\{i(b_{1}u_{1}+b_{2}u_{2})-\frac{1}{2}\sum_{j,k=1,2}\Sigma_{jk}u_{j}u_{k}\right.\right.$
$+\left.\left.\int_{\mathbb{R}^{2}\setminus\{0\}}\left(e^{i(u_{1}x+u_{2}y)}-1-i(u_{1}x+u_{2}y)I_{(0,1]}(x^{2}+y^{2})\right)\,\nu(\mathrm{d}x,\mathrm{d}y)\right\}\right).$ (7)
Here, $\Sigma=(\Sigma_{jk})_{j,k=1,2}$ is the covariance matrix of the Brownian motion in the Lévy-Itô decomposition, and $\nu(\mathrm{d}x,\mathrm{d}y)$ is the Lévy measure on $\mathbb{R}^{2}\setminus\{0\}$ satisfying
$\int_{\mathbb{R}^{2}\setminus\{0\}}\min(1,x^{2}+y^{2})\,\nu(\mathrm{d}x,\mathrm{d}y)<\infty.$
@@
Let $(X_{t},Y_{t})_{t\geq 0}$ be a two dimensional Lévy process with the Lévy-Khinchine formula (1). Define
$\rho=\left\{\begin{array}[]{ll}\frac{\Sigma_{12}}{\sqrt{\Sigma_{11}\Sigma_{22}}},&\Sigma_{11}\Sigma_{22}>0,\\ 0,&\Sigma_{11}\Sigma_{22}=0.\end{array}\right.$
@@
holds for all measurable functions $\varphi(x)$ and $\psi(y)$ such that $\varphi(0)=\psi(0)=0$. Then, we have that
@@
Next, we consider applications of Theorem 1.2 to stable random vectors. Lancaster have obtained an expression of the maximal correlation coefficients for jointly Gaussian random variables, see (3). It is natural to ask for an expression for jointly stable random variables. The index $\alpha$ of stability takes values in $(0,2]$. When $\alpha=2$, it must be jointly Gaussian. Therefore, we consider the case $\alpha\in(0,2)$. Suppose $(X,Y)$ is a two-dimensional stable random vector. Then, it is infinitely divisible and has the following Lévy-Khinchine formula:
$E[e^{i(u_{1}X+u_{2}Y)}$ $]=\exp\Big{(}\Big{\{}i(b_{1}u_{1}+b_{2}u_{2})$
$+\left.\left.\int_{\mathbb{R}^{2}\setminus\{0\}}\left(e^{i(u_{1}x+u_{2}y)}-1-i(u_{1}x+u_{2}y)I_{(0,1]}(x^{2}+y^{2})\right)\,\nu(\mathrm{d}x,\mathrm{d}y)\right\}\right).$ (8)
Here, the Lévy measure $\nu$ takes the form
$\nu(B)=\int_{0}^{\infty}\int_{\mathbb{R}/(2\pi\mathbb{Z})}I_{B}(r\cos\theta,r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta),\quad B\in\mathcal{B}(\mathbb{R}^{2}\setminus\{0\}),$ (9)
where the spectral measure $\tau(\mathrm{d}\theta)$ is a finite Borel measure.
@@
Let $(X,Y)\in\mathbb{R}^{2}$ be a stable random vector with the index $\alpha\in(0,2)$. Let
and
Then, we have that
$R(X,Y)=\mathrm{Op}(\nu)=\left\|\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}\right\|_{2}$
with the convention that $/0=0$ 0, where $\|\cdot\|_{2}$ means the spectral norm.
For random vectors $(\widetilde{X},\widetilde{Y})$ in the domain of contraction of stable laws, Theorem 1.3 provides a lower bound for $R(\widetilde{X},\widetilde{Y})$, see Remark 4.1 below.
Finally, we consider analogs and generalizations of Dembo-Kagan-Shepp inequality. We start by the following proposition, which is proved by using Csáki-Fischer identity.
Proposition 1.4.
Let $(X_{n})_{n\geq 1}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $(B_{n})_{n\geq 1}$ be i.i.d. Bernoulli random variables such that $P(B_{n}=1)=1-P(B_{n}=0)=p_{n}$ and $(B_{n})_{n\geq 1}$ is independent of $(X_{n})_{n\geq 1}$. Define
$Y_{n}=\left\{\begin{array}[]{ll}X_{n},&\text{if }B_{n}=1,\\ \partial,&\text{if }B_{n}=0.\end{array}\right.$
Then, for $n\geq 1$, we have that
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))=\sqrt{\max(p_{1},p_{2},\ldots,p_{n})}.$
Let $T=\{i=1,2,\ldots,n:B_{n}=1\}$. Then, $T$ is a random subset of $[n]:=\{1,2,\ldots,n\}$. Note that $\max(p_{1},p_{2},\ldots,p_{n})=\max\{i\in[n]:P(i\in T)\}$. Hence, the following theorem due to Madiman and Barron is a generalization of Proposition 1.4.
@@
Let $X_{1},X_{2},\ldots,X_{n}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $T$ be a random subset of $[n]=\{1,2,\ldots,n\}$. Suppose that $T$ is independent from $(X_{1},X_{2},\ldots,X_{n})$. For $i=1,2,\ldots,n$, define
@@
Then, for $n\geq 1$, we have that
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))=\sqrt{\max\{i\in[n]:P(i\in T)\}}.$
The lower bound $R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\geq\sqrt{\max\{i\in[n]:P(i\in T)\}}$ is a simple consequence of the fact $R(X_{i},Y_{i})=\sqrt{P(i\in T)}$, which follows from Proposition 1.4 for $n=1$. The most interesting part is the upper bound
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{\max\{i\in[n]:P(i\in T)\}}.$
This upper bound is known in a different form by Madiman and Barron, see [MB07, Lemmas 2 and 4]. In the same paper, this upper bound is further used to deduce generalized entropy power inequalities, monotonicity of Fisher information and the Fisher information inequality of Stam. It is important follow-up work of the break-through work [ABBN04] by Artstein, Ball, Barthe and Naor on the monotonicity of Shannon entropy.
@@
The random variables $X_{1},X_{2},\ldots,X_{n}$ are not necessarily identically distributed.
@@
The Madiman-Barron inequality is a generalization of Dembo-Kagan-Shepp inequality. Indeed, let $X_{1},X_{2},\ldots,X_{n}$ be independent random variables taking values in a general measurable space $(F,\mathcal{F})$. For fixed $m\in[n]=\{1,2,\ldots,n\}$, let $T$ be a uniform subset of $\{1,2,\ldots,n\}$ with size $m$. For $i=1,2,\ldots,n$, define
where $\partial$ is some special point in $(F,\mathcal{F})$. Then, by the Madiman-Barron inequality, for $n\geq 1$,
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{m/n}.$
Suppose $F$ is a vector space. Then, $(\sum_{i=1}^{n}X_{n},\sum_{i=1}^{n}Y_{n}1_{Y_{n}\neq\partial})$ has the same joint distribution as $(S_{n},S_{m})$ in (4). Hence, we have that
Note that Theorem 1.5 gives the maximal correlation coefficient between a finite independent sequence and its random sub-sequence. Hence, it is natural to ask for a generalization for two random sub-sequences of the same finite independent sequence. We provide an answer by the following theorem:
@@
Let $X_{1},X_{2},\ldots,X_{n}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $S$ and $T$ be two random subsets of $[n]=\{1,2,\ldots,n\}$. Suppose that $(S,T)$ is independent from $(X_{1},X_{2},\ldots,X_{n})$. For $i=1,2,\ldots,n$, define
@@
Then, for $n\geq 1$, we have that
@@
$\sum_{s,t:j\in s\cap t}P(S=s,T=t)\alpha_{s}\beta_{t}\leq r\sqrt{\sum_{s:j\in s}P(S=s)\alpha_{s}^{2}}\sqrt{\sum_{t:j\in t}P(T=t)\beta_{t}^{2}}$ (11)
for arbitrary real constants $\alpha_{s}$ and $\beta_{t}$. In particular, when $S$ and $T$ are independent, we have $R(S,T)=0$ and $r_{j}=\sqrt{P(j\in S)P(j\in T)}$.
@@
According to Theorem 1.6, in order to calculate $R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))$, we need to know $R(S,T)$. When $S$ and $T$ are dependent, we have no answer in general. However, we have a partial answer in the following special case.
@@
Let $T$ be a uniform subset of $[n]=\{1,2,\ldots,n\}$ of size $m$. Given $T$, the random set $S$ is a uniform subset of $T$ of size $k\leq m$. Then, we have that
@@
with the convention that $/0=0$ 0.
Recall that Yu [Yu08, Theorem 4.1] proved (5). We give a generalization of the upper bound $R\left(\sum_{i=1}^{m}X_{i},\sum_{j=\ell+1}^{n}X_{j}\right)\leq\frac{m-\ell}{\sqrt{m(n-\ell)}}$ in (5) as an application of Theorem 1.6:
Corollary 1.8.
We use the notation and assumptions in Theorem 1.6. Assume that $(S,T)$ is uniformly distributed in the following set $\{(s,t):s,t\subset[n],|s|=k,|t|=m,|s\cap t|=\ell\}$. Then, we have that
$R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))=\frac{\ell}{\sqrt{mk}}.$
By adapting the arguments in [Cou16] and [MB07], we find that Theorem 1.6 implies the following results in information theory:

### v3

@@ Abstract
We study the maximal correlation coefficient $R(X,Y)$ between two stochastic processes $X$ and $Y$. In the case when $(X,Y)$ is a random walk, we find $R(X,Y)$ using the Csáki-Fischer identity and the lower semicontinuity of the map $\text{Law}(X,Y)\to R(X,Y)$. When $(X,Y)$ is a two-dimensional Lévy process, we express $R(X,Y)$ in terms of the Lévy measure of the process and the covariance matrix of the diffusion part of the process. Consequently, for a two-dimensional $\alpha$-stable random vector $(X,Y)$ with $0<\alpha<2$, we express $R(X,Y)$ in terms of $\alpha$ and the spectral measure $\tau$ of the $\alpha$-stable distribution. We also establish analogs and extensions of the Dembo-Kagan-Shepp-Yu inequality and the Madiman-Barron inequality.
@@ 1 Introduction
For two square-integrable non-degenerate real-valued random variables $X$ and $Y$, the Pearson correlation coefficient $\rho(X,Y)$ is defined by
@@
The correlation coefficient $\rho(X,Y)$ measures the linear dependence of $X$ and $Y$. It is well known that $\rho(X,Y)=0$ if $X$ and $Y$ are independent and square-integrable. However, the converse is not true in general. For two non-degenerate random variables $X$ and $Y$, the maximal correlation coefficient between $X$ and $Y$, introduced by Gebelein [Geb41], is given by
@@
where the supremum is over all measurable functions $\varphi$ and $\psi$ such that
$0<E[(\varphi(X))^{2}]<\infty,\quad 0<E[(\psi(Y))^{2}]<\infty.$
If $X$ or $Y$ is degenerate, we set $R(X,Y)=0$. The quantity $R(X,Y)$ measures the dependence between $X$ and $Y$, with $R(X,Y)=0$ indicating independence, see [Sar58].
According to [R5́9, Eq. (29)], the maximal correlation coefficient has the following alternative definition: For a random variable $X$, let $L^{2}_{0}(X)$ denote the Hilbert space of square-integrable mean 0 real-valued random variables $Z$ that are measurable with respect to the sigma-field $\sigma(X)$ generated by $X$. For two random variables $X$ and $Y$, the maximal correlation coefficient $R(X,Y)$ is equal to the operator norm of the conditional expectation $\varphi(X)\mapsto E(\varphi(X)|Y)$ from $L^{2}_{0}(X)$ to $L^{2}_{0}(Y)$. Equivalently, we have the following expression:
$R^{2}(X,Y)=\sup\{E[(E[\varphi(X)|Y])^{2}]:E[\varphi(X)]=0,E[\varphi^{2}(X)]=1\}.$ (2)
Recently, the maximal correlation coefficient has been generalized by Dadoun and Youssef [DY21] to the context of free probability. However, in this paper, we will concentrate on the classical notion of maximal correlation.
The maximal correlation coefficient plays an important role in various areas of probability and statistics, including information theory [Cou16, MB07], the hyper-contractivity ribbon of a pair of random variables and the impossibility of non-interactive simulations of joint distributions [AG76, KA12, KA16], the optimal transformation for regression [BF85], the spectral gap of Markov chains and the convergence theory of data augmentation algorithms [LWK94].
As noted by Rényi in [R5́9], it is often difficult to find the exact value of the maximal correlation coefficient, although it may not seem so at first sight. One reason is that the supremum in (1) generally cannot be replaced by a maximum. The explicit value of the maximal correlation coefficient $R(X,Y)$ is only known in a few cases. For example, if $(X,Y)$ is jointly Gaussian, then
@@
see [CF60, Example 5 in Section 2]. The exact value of $R(X,Y)$ was also found in the case when $(X,Y)$ follows a multinomial distribution or multivariate hypergeometric distribution in [CF63, Section 6]. Bücher and Staud [BS25] obtained the maximal correlation coefficient for the bivariate Marshall-Olkin exponential distribution. There are also several results for maximal correlation coefficients by using orthogonal polynomials, as discussed in [LBCnM06, LBSMn14, LBSMn98, Nev92, PX13, SM85, Ter83]. Among these results, [SM85] gives the maximal correlation coefficients for Dirichlet distributions, which provides sharp upper bounds for the maximal correlation coefficients for order statistics studied in [SM85, Ter83]. In [LBSMn98, Nev92], they studied the maximal correlation between the $i$-th and $j$-th records. In [PX13], they provided a unified approach to obtain the maximal correlation coefficient for a subclass of Lancaster distributions. There are also some results on the maximum correlation coefficients for arbitrary distributions. For instance, if $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are independent, then
@@
This is known as the Csáki-Fischer identity [CF63, Theorem 6.2], see also [Wit75, Theorem 1]. If $X_{1},X_{2},\ldots,X_{n}$ are independent and identically distributed (i.i.d.) non-degenerate real-valued random variables, then for $m\leq n$, we have
@@
where $S_{k}=\sum_{i=1}^{k}X_{i}$ is the partial sum of $(X_{i})_{i}$, see [BDK05, DKS01, Nov04]. The upper bound
$R(S_{n},S_{m})\leq\sqrt{m/n}$
is known as the Dembo-Kagan-Shepp (DKS) inequality. Yu [Yu08, Theorem 4.1] further generalized (4) to
@@
In this paper, we are interested in the maximal correlation coefficient between two stochastic processes, e.g., two random walks, two Lévy processes, or the maximal correlation coefficient between two randomly chosen sub-vectors of a third common random vector. Our main results are as follows:
1.1 Maximal correlation coefficients of random walks
Let $(\xi_{n},\eta_{n})_{n\geq 1}$ be i.i.d. random vectors. Let $(S_{n},T_{n})_{n\geq 0}$ be a random walk starting from 0 with increments $(\xi_{n},\eta_{n})$, that is,
@@ 1.1 Maximal correlation coefficients of random walks
Let $S=(S_{n})_{n\geq 0}$ and $T=(T_{n})_{n\geq 0}$. Our first main result is
Theorem 1.1.
For any $m\geq 1$,
@@
The first equality is a direct consequence of the Csáki-Fischer identity (Theorem 2.2) since
@@
Our main contribution is the proof of the second equality, where we use the lower semicontinuity of $\text{Law}(X,Y)\mapsto R(X,Y)$ (Lemma 2.3). For self-containedness, we also provide a new probabilistic proof of the Csáki-Fischer identity. As a consequence of Theorem 1.1, the central limit theorem and the lower semicontinuity of $\text{Law}(X,Y)\mapsto R(X,Y)$ (Lemma 2.3), we provide a new proof of (3) in Subsection 3.2. This is the main motivation for studying the maximal correlation coefficients for random walks. Similarly, using the law of rare events instead of the central limit theorem, we see that (3) holds for bivariate Poisson distributions.
1.2 Maximal correlation coefficients of two-dimensional Lévy processes and bivariate stable distributions
By analyzing maximal correlation coefficients for random walks, we have established the maximal correlation coefficients for bivariate Gaussian vectors. Consequently, it seems logical to investigate the problem of bivariate stable distributions through the analysis of Lévy processes. We refer to Applebaum’s book [App09] for properties of Lévy processes.
@@ 1.2 Maximal correlation coefficients of two-dimensional Lévy processes and bivariate stable distributions
Let $(X_{t},Y_{t})$ be a two-dimensional Lévy process with the characteristic triple $(b,\Sigma,\nu)$. Define
$\rho=\left\{\begin{array}[]{ll}\Sigma_{12}/\sqrt{\Sigma_{11}\Sigma_{22}},&\Sigma_{11}\Sigma_{22}>0,\\ 0,&\Sigma_{11}\Sigma_{22}=0.\end{array}\right$.
@@
holds for all measurable functions $\varphi(x)$ and $\psi(y)$ such that $\varphi(0)=\psi(0)=0$. Then we have that
@@
As an application of the above result, we can get the maximal correlation coefficient of a bivariate stable distribution with the stability index $\alpha\in(0,2)$.
@@
Let $(X,Y)\in\mathbb{R}^{2}$ be a stable random vector with index $\alpha\in(0,2)$. Let $_{++} ,C+C_{+-}$, $_{-+}, CC_{--}$, $D^{x}_{+}$, $D^{x}_{-}$, $D^{y}_{+}$, $D^{y}_{-}$ be some integrals defined in Lemma 4.4. Then we have that
$R(X,Y)=\left\|\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&\quad C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&\quad C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}\right\|_{2}$
with the convention that $0/0=0$, where $\|\cdot\|_{2}$ denotes the spectral norm.
[…]
1.3 Extensions of the Dembo-Kagan-Shepp-Yu inequality
The Dembo-Kagan-Shepp-Yu (DKSY) inequality
$R\left(\sum_{i=1}^{m}X_{i},\sum_{j=\ell+1}^{n}X_{j}\right)\leq\frac{m-\ell}{\sqrt{m(n-\ell)}}$ (7)
is a sharp inequality for two partial sums of i.i.d. non-degenerate real-valued random variables. It is also natural to consider other functions instead of partial sums. Therefore, we are interested in the maximal correlation coefficient between two randomly chosen subvectors of a common third random vector. Madiman and Barron [MB07] found a sharp upper bound for the maximal correlation coefficient between a random vector and its randomly chosen subvector.
@@ 1.3 Extensions of the Dembo-Kagan-Shepp-Yu inequality
Let $X_{1},X_{2},\ldots,X_{n}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $T$ be a random subset of $n]=\{1,2,\ldots,n\}$. Suppose that $T$ is independent of $(X_{1},X_{2},\ldots,X_{n})$. For $i=1,2,\ldots,n$, define
@@
Then for $n\geq 1$, we have
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))=\sqrt{\max\{P(i\in T):i\in[n]\}}.$
@@
The random variables $X_{1},X_{2},\ldots,X_{n}$ need not be identically distributed.
@@
The Madiman-Barron inequality is a generalization of the Dembo-Kagan-Shepp inequality, see Appendix C.
Madiman-Barron [MB07] proved the following Madiman-Barron inequality
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{\max\{P(i\in T):i\in[n]\}}.$
This upper bound appears in a different form in [MB07, Lemmas 2 and 4]. In the same paper, this upper bound is further used to deduce generalized entropy power inequalities, the monotonicity of Fisher information, and the Fisher information inequality of Stam. It is an important follow-up work of the breakthrough work [ABBN04] by Artstein, Ball, Barthe and Naor on the monotonicity of the Shannon entropy. However, we find no proof in the literature for the lower bound
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\geq\sqrt{\max\{P(i\in T):i\in[n]\}}.$
We will give a proof in Subsection 5.1. For the lower bound, it is crucial that $\partial$ is outside of $F$, see Remark 5.1. However, for the upper bound, $\partial$ is not necessarily outside of $F$.
Next, we extend Madiman-Barron’s result to the case of the maximal correlation coefficient of two randomly chosen subvectors:
@@
Let $X_{1},X_{2},\ldots,X_{n}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $S$ and $T$ be two random subsets of $n]=\{1,2,\ldots,n\}$. Suppose that $(S,T)$ is independent of $(X_{1},X_{2},\ldots,X_{n})$. For $i=1,2,\ldots,n$, define
@@
Then for $n\geq 1$, we have that
@@
$\sum_{s,t:j\in s\cap t}P(S=s,T=t)\alpha_{s}\beta_{t}\leq r\sqrt{\sum_{s:j\in s}P(S=s)\alpha_{s}^{2}}\sqrt{\sum_{t:j\in t}P(T=t)\beta_{t}^{2}},\quad\forall\text{ real }\alpha_{s},\beta_{t}.$ (9)
In particular, when $S$ and $T$ are independent, we have $R(S,T)=0$ and $r_{j}=\sqrt{P(j\in S)P(j\in T)}$.
@@
According to Theorem 1.5, in order to calculate $R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))$, we need to know $R(S,T)$. When $S$ and $T$ depend on each other, we have no good control over $R(S,T)$. However, we have a partial answer in the following special case:
@@
Let $T$ be a uniform subset of $n]=\{1,2,\ldots,n\}$ of size $m$. Given $T$, let $S$ be a uniform subset of $T$ with size $k\leq m$. Then we have that
@@
with the convention that $0/0=0$.
As a corollary, we extend the Dembo-Kagan-Shepp-Yu inequality:
Corollary 1.7 (Dembo-Kagan-Shepp-Yu).
Let us consider i.i.d. non-degenerate random variables $X_{1},X_{2},\ldots,X_{n}$ taking values in a general measurable space $(F,\mathcal{F})$. Then for $1\leq\ell+1\leq m\leq n$, we have that
$R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{j=\ell+1}^{n}\delta_{X_{j}}\right)=\frac{m-\ell}{\sqrt{m(n-\ell)}},$ (10)
where $\delta_{x}$ is the Dirac measure at $x$.
Remark 1.3.
The original version of the Dembo-Kagan-Shepp-Yu inequality (7) is stated for real-valued random variables. We believe that a similar result also holds for random vectors by adapting their arguments. The upper bound
$R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{j=\ell+1}^{n}\delta_{X_{j}}\right)\leq\frac{m-\ell}{\sqrt{m(n-\ell)}}$
in Corollary 1.7 is equivalent to the generalized version of DKSY inequality for random vectors. However, the generalized version for random vectors is not clearly written in the literature. So, we decide to state Corollary 1.7 and provide a proof for self-containedness. In the present paper, Corollary 1.7 is used to deduce Proposition 5.5.
Remark 1.4.
In general, the equation (10) does not hold for independent non-degenerate random variables $X_{1},X_{2},\ldots,X_{n}$ with different distributions. Indeed, for some particular distributions, there is a one-to-one correspondence between $\sum_{i=a}^{b}\delta_{X_{i}}$ and $(X_{a},X_{a+1},\ldots,X_{b})$ for all $1\leq a\leq b\leq n$. In such cases, we have that
$R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{j=\ell+1}^{n}\delta_{X_{j}}\right)\geq R(X_{\ell+1},X_{\ell+1})=1.$
At present, we do not have a satisfactory extension of the Dembo-Kagan-Shepp-Yu inequality for random variables with different distributions. However, we have such an extension for the Dembo-Kagan-Shepp inequality, see Lemma 5.3.
By applying Theorem 1.5 and adapting the arguments in [Cou16, MB07], we can get the following result in information theory: